# Resolution and Robustness Bounds for Reconstructive Spectrometers

### v1

@@ Abstract
In this Letter, we develop a broadly-applicable figure of merit for the performance of a reconstructive spectrometer, and analyze its implications for the device’s desired physical characteristics. We first abstract away most implementation details, retaining only the following assumptions: (i) the spectrum is inferred from light intensity measurements on a fixed set of output channels; (ii) each intensity measurement is subject to independent Gaussian noise; and (iii) there are no additional relevant details, such as priors about the spectrum (e.g., sparsity). The number of measurement channels may be smaller than the number of frequency channels, so that the inference problem is under-determined. Under these conditions, the theory of Fisher information [frieden1998physics] implies that the noise-induced reconstruction error is bounded below by $\sigma_{\epsilon}^{2}\operatorname{Tr}[(\mathbf{A}^{T}\mathbf{A})^{+}]$, where $\sigma_{\epsilon}$ is the noise level, $\mathbf{A}$ is the normalized spectral transmission matrix, and $(\cdots)^{+}$ denotes the Moore–Penrose pseudoinverse. By leveraging on empirically-grounded features of generic $\mathbf{A}$ matrices [fyodorov2016random, lehmann1995chaotic, fyodorov1997statistics, popoff2010measuring, goetschy2013filtering, cao2015dielectric], we derive expressions for the noise-induced reconstruction error and effective spectral resolution, in terms of $\Gamma_{\mathrm{corr}}$, the average transmittance, and the number of frequency and measurement channels. In particular, our results clarify the necessary conditions for a reconstructive spectrometer to achieve “super-resolution” below the scale set by $\Gamma_{\mathrm{corr}}$.
We subject the theory to two sets of numerical tests: a random matrix theory (RMT) model, and full-wave simulations. Over a wide range of conditions, the theoretical predictions closely match the actual error rates, either using the pseudoinverse or a neural network for inference. We also demonstrate that the super-resolution regime is achievable in both kinds of models. In the full-wave simulations, we show that combining our results with the scaling laws for diffusive transport [akkermans2007mesoscopic] gives an accurate prediction for the optimal device size, without requiring a brute-force search. Our theory thus yields useful insights to guide the design of reconstructive spectrometers.
@@
$\Delta\omega_{\min}\approx\frac{\sqrt{2}M}{N}\frac{\pi\Gamma_{\text{corr}}}{\ln\mathcal{R}_{\text{SNR}}},$ (16)
@@
Next, we analyze how the key scattering properties vary with $L$. As shown in Fig. 3(c), both the mean transmittance $T_{0}$ and the correlation length $\Gamma_{\mathrm{corr}}$ decrease with increasing $L$, consistent with the theoretical scaling laws for diffusive transport: $T_{0}\propto\ell/L$ and $\Gamma_{\mathrm{corr}}\propto D/L^{2}$, where $\ell$ is the transport mean free path and $D$ is the diffusion constant [akkermans2007mesoscopic]. Hence, we can deduce the optimal cavity size without conducting a brute-force search. The reconstruction error $\mathcal{E}$ is minimized by optimizing $\mathcal{E}[a(L)]\propto\tilde{J}[a(L)]/a(L)$, where $a(L)$ is the normalized correlation parameter. Solving $d\mathcal{E}/da=0$ numerically for $M=9$ and $N=100$, using Eq. (14), yields $a_{\mathrm{opt}}\approx 7.46$. Using the fitted diffusion constant $D\approx 120\,\mu\textrm{m}^{2}/\mathrm{THz}$ and frequency channel spacing $\Delta\omega=0.0439$ THz, we obtain an optimal cavity size of $L_{\mathrm{opt}}=\sqrt{D/(a_{\mathrm{opt}}\cdot\Delta\omega)}\approx 17.9\,\mu\text{m}$. This matches the actual minimum of $\operatorname{Tr}[\mathbf{G}^{+}]$, as shown in Fig. 3(d). The plot also illustrates the key trade-off highlighted by the theory: for $L<L_{\mathrm{opt}}$, performance is limited by excessive spectral correlations, while for $L>L_{\mathrm{opt}}$, it is degraded by low optical throughput.
@@
In Fig. 3(e), we plot the effective resolution $\Delta\omega_{\textrm{min}}$ and $\Gamma_{\text{corr}}$ versus $L$. It can be seen that the spectrometer operates in the super-resolution regime below a certain cavity size, and the previously-obtained value of $L_{\mathrm{opt}}$ lies in this regime. Note that our previously-derived $L_{\mathrm{opt}}$ is near, but not equal to, the minimum of $\Delta\omega_{\textrm{min}}$ in this plot. This is because these are two related but distinct design targets that one has to choose between: the former maximizes the spectrometer’s overall fidelity (by minimizing the noise-induced MSE), while the latter optimizes the effective frequency resolution at a given noise level and detection threshold. To clarify this, Fig. 3(f) shows the effective resolution $\Delta\omega_{\textrm{min}}$ versus detection threshold $\delta_{\mathrm{th}}^{2}$ for two cavity sizes: (i) $L=L_{\mathrm{opt}}$, and (ii) $L=25\,\mu\textrm{m}$, where the resolution curve in Fig. 3(e) reaches its minimum. The latter indeed yields a finer spectral resolution when operating at $\delta_{\mathrm{th}}^{2}=2\times 10^{-4}$, which is the detection threshold used in Fig. 3(e); however, for smaller values of $\delta_{\mathrm{th}}^{2}$, it yields a worse (i.e., larger) resolution.
@@
Conclusions.—The performance of a reconstructive spectrometer is largely determined by the variance bound $\operatorname{Tr}[\mathbf{G}^{+}]$, derived from its spectral transmission matrix. In the regime where noise is present but not dominant, this quantity accurately describes the noise-induced error for not only pseudoinverse-based reconstruction, but also nonlinear inference using a neural network, implying that this criterion is reasonably algorithm-independent. We have derived an expression for $\operatorname{Tr}[\mathbf{G}^{+}]$, and hence the effective frequency resolution, in terms of physically meaningful variables: the spectral correlation length $\Gamma_{\mathrm{corr}}$, mean transmittance $T_{0}$, and the number of frequency and intensity measurement channels. Our theory gives successful predictions for both RMT models and an exemplary full-wave simulation of an on-chip device. These tests confirm that $\Gamma_{\mathrm{corr}}$ is not the sole determinant of performance for a reconstructive spectrometer; $T_{0}$ and other variables also play a role and can lead to trade-offs.
One noteworthy limitation of our analysis is that it assumes the spectrometer follows ordinary scattering properties, including conventional speckle statistics with Lorentzian spectral correlations [ericson1960fluctuations, lehmann1995chaotic, fyodorov1997statistics, weidenmuller2009random]. These assumptions may be violated by individual extremal scatterers. We have probed this issue by conducting numerical inverse design to minimize $\operatorname{Tr}[\mathbf{G}^{+}]$ [Molesky2018, Kroker2024, yu2025wavelength]. As detailed in the Supplemental Materials, the optimizer initially matches our theory (i.e., changing $T_{0}$ and $\Gamma_{\mathrm{corr}}$ as the theory predicts), but eventually achieves an even lower value of $\operatorname{Tr}[\mathbf{G}^{+}]$ by generating structures with highly non-Lorentzian correlations [SM]. In future, it would be interesting to explore whether the extremal statistics of RMT models [Lakshminarayan2008] could describe such optimized structures, and thereby arrive at an explanatory theoretical framework for the ultimate limits of reconstructive spectrometers.
@@
We thank A. Nussupbekov, B. Zhu, and S. G. Johnson for helpful discussions. This work was supported by the Singapore National Research Foundation (NRF) under the NRF Investigatorship NRF-NRFI08-2022-0001, and Competitive Research Program (CRP) Nos. NRF-CRP23-2019-0005, NRF-CRP23-2019-0007, and NRF-CRP29-2022-0003.
@@
As noted in the conclusion of our paper, it may be possible for individual extremal scatterers to surpass the performance bounds predicted by our theory, due to certain assumptions embedded in the theory. In particular, we have assumed that the scatterer obeys conventional speckle statistics with Lorentzian spectral correlations—which holds empirically for many real disordered scatterers and can be derived from RMT [ericson1960fluctuations, lehmann1995chaotic, fyodorov1997statistics, weidenmuller2009random], but is not an immutable physical law.

### v2

@@ Abstract
In this Letter, we develop a broadly-applicable figure of merit for the performance of a reconstructive spectrometer, and analyze its implications for the device’s desired physical characteristics. We first abstract away most implementation details, retaining only the following assumptions: (i) the spectrum is inferred from light intensity measurements on a fixed set of output channels; (ii) each intensity measurement is subject to independent Gaussian noise; and (iii) there are no additional relevant priors about the spectrum (e.g., sparsity). The number of measurement channels may be smaller than the number of frequency channels, so that the inference problem is under-determined. Under these conditions, the theory of Fisher information [frieden1998physics] implies that the noise-induced reconstruction error is bounded below by $\sigma_{\epsilon}^{2}\operatorname{Tr}[(\mathbf{A}^{T}\mathbf{A})^{+}]$, where $\sigma_{\epsilon}$ is the noise level, $\mathbf{A}$ is the normalized spectral transmission matrix, and $(\cdots)^{+}$ denotes the Moore–Penrose pseudoinverse. By leveraging on empirically-grounded features of generic $\mathbf{A}$ matrices [fyodorov2016random, lehmann1995chaotic, fyodorov1997statistics, popoff2010measuring, goetschy2013filtering, cao2015dielectric], we derive expressions for the noise-induced reconstruction error and effective resolution in terms of $\Gamma_{\mathrm{corr}}$, the average transmittance, and the number of frequency and measurement channels. In particular, our results clarify the necessary conditions for a reconstructive spectrometer to achieve “super-resolution” below the scale set by $\Gamma_{\mathrm{corr}}$.
We subject the theory to two sets of numerical tests: a random matrix theory (RMT) model, and full-wave simulations. Over a wide range of conditions, the theoretical predictions closely match the actual error rates, either using the pseudoinverse or a neural network for inference. We also demonstrate that the super-resolution regime is achievable in both kinds of models. In the full-wave simulations, we show that combining our results with the scaling laws for diffusive transport [akkermans2007mesoscopic] predicts an optimal device size without a brute-force search. Our theory thus yields useful insights to guide the design of reconstructive spectrometers.
@@
$\Delta\omega_{\min}\approx\frac{\sqrt{2}M}{N}\frac{\pi\Gamma_{\text{corr}}}{\ln\mathcal{R}_{\text{SNR}}}.$ (16)
@@
Figure 3(c) shows how the scattering properties depend on $L$. Both $T_{0}$ and $\Gamma_{\mathrm{corr}}$ vary according to the scaling laws for diffusive transport: $T_{0}\propto\ell/L$ and $\Gamma_{\mathrm{corr}}\propto D/L^{2}$, where $\ell$ is the transport mean free path and $D$ is the diffusion constant [akkermans2007mesoscopic]. Using this, we can deduce the optimal cavity size without a brute-force search. We aim to minimize the reconstruction error $\mathcal{E}[a(L)]\propto\tilde{J}[a(L)]/a(L)$; solving $d\mathcal{E}/da=0$ numerically for $M=9$ and $N=100$, using Eq. (14), yields $a_{\mathrm{opt}}\approx 7.46$. Using the fitted diffusion constant $D\approx 120\,\mu\textrm{m}^{2}/\mathrm{THz}$ and frequency channel spacing $\Delta\omega=0.0439$ THz, we obtain an optimal cavity size $L_{\mathrm{opt}}=\sqrt{D/(a_{\mathrm{opt}}\cdot\Delta\omega)}\approx 17.9\,\mu\text{m}$, matching the observed minimum of $\operatorname{Tr}[\mathbf{G}^{+}]$ as shown in Fig. 3(d). This also illustrates a key trade-off highlighted by the theory: for $L<L_{\mathrm{opt}}$, performance is limited by excessive spectral correlations, whereas for $L>L_{\mathrm{opt}}$ it is limited by low optical throughput.
@@
In Fig. 3(e), we plot the effective resolution $\Delta\omega_{\textrm{min}}$ and $\Gamma_{\text{corr}}$ versus $L$. It can be seen that the spectrometer operates in the super-resolution regime below a certain cavity size, and the previously-obtained value of $L_{\mathrm{opt}}$ lies in this regime. Note that our previously-derived $L_{\mathrm{opt}}$ is near, but not equal to, the minimum of $\Delta\omega_{\textrm{min}}$ in this plot. These are two related but distinct design targets one has to choose between: the former maximizes the overall fidelity (by minimizing the noise-induced MSE), while the latter optimizes the effective frequency resolution at a given noise level and detection threshold. To clarify this, Fig. 3(f) shows the effective resolution $\Delta\omega_{\textrm{min}}$ versus detection threshold $\delta_{\mathrm{th}}^{2}$ for two cavity sizes: (i) $L=L_{\mathrm{opt}}$, and (ii) $L=25\,\mu\textrm{m}$, where the resolution curve in Fig. 3(e) reaches its minimum. The latter indeed yields a finer spectral resolution when operating at $\delta_{\mathrm{th}}^{2}=2\times 10^{-4}$, the detection threshold used in Fig. 3(e); for smaller $\delta_{\mathrm{th}}^{2}$, it gives a worse (i.e., larger) resolution.
@@
Conclusions.—The performance of a reconstructive spectrometer is largely determined by the variance bound $\operatorname{Tr}[\mathbf{G}^{+}]$, derived from its spectral transmission matrix. When noise is present but non-dominant, this quantity accurately describes the noise-induced error for not only pseudoinverse-based reconstruction, but also inference using a neural network, implying that the criterion is reasonably algorithm-independent. We have derived expressions for $\operatorname{Tr}[\mathbf{G}^{+}]$, and hence the effective frequency resolution, in terms of physically meaningful variables: the spectral correlation length $\Gamma_{\mathrm{corr}}$, mean transmittance $T_{0}$, the number of intensity measurement channels, etc. The theory yields successful predictions for both RMT models and a full-wave simulation of an on-chip device. The tests confirm that $\Gamma_{\mathrm{corr}}$ is not the sole determinant of performance for a reconstructive spectrometer; notably, $T_{0}$ also plays a role and can lead to a trade-off.
A key assumption we have made is that the scatterer exhibits conventional speckle statistics with Lorentzian spectral correlations [ericson1960fluctuations, lehmann1995chaotic, fyodorov1997statistics, weidenmuller2009random]. This assumption holds for almost all scattering-based reconstructive spectrometers implemented to date, including those based on photonic chips [redding2013compact], multimode fibers [redding2014using], colloids [bao2015colloidal], and nanostructures [yang2019single, Lee2025Reconstructive]—but it is not a universal law. In the Supplemental Materials [SM], we demonstrate that when a scatterer is expressly inverse-designed [Molesky2018, Kroker2024] to minimize $\operatorname{Tr}[\mathbf{G}^{+}]$ [yu2025wavelength], the generated structure eventually acquires highly non-Lorentzian correlations while also surpassing the performance predicted by our theory. This hints that non-Lorentzian speckle statistics, which have previously been studied in the context of fundamental wave physics [Lakshminarayan2008, chabanov2000statistical], may have an important application in this area. In future, it will be desirable to find additional theoretical techiques to model such systems, such as the extremal statistics of RMT models [Lakshminarayan2008], which would allow us to characterize the ultimate limits of reconstructive spectrometry.
@@
We thank A. Nussupbekov, B. Zhu, F. J. N. Jorgensen, and S. G. Johnson for helpful discussions. This work was supported by the Singapore National Research Foundation (NRF) under the NRF Investigatorship NRF-NRFI08-2022-0001, and Competitive Research Program (CRP) Nos. NRF-CRP23-2019-0005, NRF-CRP23-2019-0007, and NRF-CRP29-2022-0003.
@@
As noted in the conclusion of our paper, it may be possible for individual extremal scatterers to surpass the performance bounds predicted by our theory, due to certain assumptions embedded in the theory. In particular, we have assumed that the scatterer obeys conventional speckle statistics with Lorentzian spectral correlations—which holds empirically for many real disordered scatterers and can be derived from RMT [ericson1960fluctuations, lehmann1995chaotic, fyodorov1997statistics, weidenmuller2009random], but is not immutable.